OpenStudy (anonymous):

Find the vector and parametric equations for the line through the point P(-2, 3, 1) and orthogonal to the plane 4x−5y+5z=−5. In vector form and parametric form (parameter t, and passing through P when t = 0):

OpenStudy (michele_laino):

Please note that the vector v=(4,-5,5) is orthogonal to your plane

OpenStudy (anonymous):

Why is that vector orthogonal to the plane?

OpenStudy (michele_laino):

so your line has to be parallel to the vector v=(4,-5,5)

OpenStudy (anonymous):

I don't understand why <4,-5,5> is orthogonal to the plane.

OpenStudy (michele_laino):

please note that the point (0,0,-1) belongs to your plane.

OpenStudy (anonymous):

I also don't understand why (0,0,-1) belongs to the plane.

OpenStudy (michele_laino):

please substitute those coordinates into the equation of your plane

OpenStudy (anonymous):

4(0) - 5(0) +5(-1) = -5, 0 - 0 - 5 = -5, -5 = -5, true statement

OpenStudy (michele_laino):

Now the generic equation of a plane passing at point P=(x,y,z) and orthogonal to vector N=(a,b,c), is: \[(X-P) \cdot N=0\] where X=x,y,z) is the generic point which belongs to the plane

OpenStudy (michele_laino):

In our case, we have N=(4,-5,5), and P=(0,0,-1), so substituting into my formula, we have: \[[(x,y,z)-(0,0,-1)] \cdot (4,-5,5)=0\] from which: \[(x,y,z+1) \cdot(4,-5,5)=0\] then: \[4x-5y+5z+5=0\] which is the equation of your plane

OpenStudy (michele_laino):

So I heve showed that your plane is orthogonal to the vector (4,-5,5)

OpenStudy (michele_laino):

oops...So I have...

OpenStudy (anonymous):

im very slow so still not getting it, takes me forever to understand anything

OpenStudy (michele_laino):

ok! Take your time, if you need please send me a message

OpenStudy (anonymous):

ok thank you :)

OpenStudy (anonymous):

closing for now so I can post another question, but will use info and post up later if I need help, ty @Michele_Laino